# Time Evidence Fusion Network: Multi-source View in Long-Term Time Series Forecasting

### v2

@@ 3.3. Basic Probability Assignment Module
To introduce non-linear characteristics into TEFN and enhance its modeling capabilities, we can modify the fuzzy membership function used in the BPA module. Instead of the linear form presented in Equation 10, we can adopt non-linear functions that capture more complex relationships within the data. Several options exist for non-linear membership functions, including:
  • Gaussian Function: This function models the data distribution as a bell-shaped curve such as Equation 3, capturing the likelihood of each data point belonging to different membership levels. The parameters of the Gaussian function can be learned to adapt to the specific characteristics of the time series data.
  • Sigmoid Function: This function smoothly transitions between 0 and 1, providing a flexible way to represent the membership of data points to different levels in Equation 11. The sigmoid function can capture both gradual and abrupt changes in membership values.
    (11) $\sigma(x)=\begin{cases}0&\text{if }x<0\\ \frac{1}{1+e^{-x}}&\text{if }x\geq 0\end{cases}$
  • Piecewise Functions: These functions consist of multiple linear or non-linear segments, allowing for more precise control over the shape of the membership function in Equation 12 where $g(x)$ is a non-linear function. This approach is useful for capturing complex patterns and relationships within the data.
    (12) $f(x)=\begin{cases}a_{1}x+b_{1}&\text{if }c_{1}x+d_{1}\leq 0\\ g_{1}(x)&\text{if }c_{2}x+d_{2}>0\text{ and }c_{3}x+d_{3}\leq 0\\ a_{2}x+b_{2}&\text{if }c_{4}x+d_{4}>0\\ \vdots&\\ a_{n}x+b_{n}&\text{otherwise}\end{cases}$
@@
3.4. Expectation Fusion
TEFN addresses the challenge of handling multivariate time series data with two distinct dimensions: the time dimension $T$ and the channel dimension $C$. This results in the paralleled generation of different mass distributions for each dimension, denoted as $m_{C}$ and $m_{T}$. To integrate these mass distributions and produce a unified prediction, TEFN employs a fusion method based on expectations, as defined in Equation 13. Here, $y_{j,k}$ represents the fusion parameters, and $m^{\prime}_{i,j,k}$ corresponds to the transformed mass function for each dimension.
(13) $\hat{y}_{Norm,D,i,j}=E_{D,F}(y_{j,k})=\sum_{\begin{subarray}{c}j\in[1,|D|]\\ k\in[1,|2^{S}|]\end{subarray}}y_{j,k}*m^{\prime}_{i,j,k}=\sum_{\begin{subarray}{c}j\in[1,|D|]\\ k\in[1,|2^{S}|]\end{subarray}}m_{i,j,k}$
The fusion operation in Equation 13 can be interpreted as a linear transformation of the mass functions, where the summation of $y_{j,k}*m^{\prime}_{i,j,k}$ effectively combines the information from different dimensions. The expectation is then calculated by summing the transformed mass functions, resulting in a single mass function representing the fused prediction.
TEFN deliberately avoids using the Dempster-Shafer Rule of Combination (DSR) for several reasons. Firstly, the computational complexity of DSR is significantly higher compared to the expectation-based fusion method. Additionally, DSR involves point-to-point multiplication of mass distributions within the same dimension, which can lead to accuracy loss when dealing with high-dimensional data due to finite precision limitations.
Secondly, DSR is sensitive to extreme distributions, particularly those with a single element having a value of 1. This sensitivity can result in the final fused result being dominated by that single distribution, potentially leading to conflicts and loss of information from other distributions.
By employing the expectation-based fusion method, TEFN achieves a balance between computational efficiency and accuracy. This approach effectively integrates information from different dimensions while mitigating the risks of conflicts and accuracy loss associated with DSR. The final normalized prediction results are then aggregated across dimensions using the equation in Equation 14 and restored to their original scale through the de-normalization layer.
(14) $\hat{y}_{norm}=\sum_{D\in\{T,C\}}\hat{y}_{Norm,D}=\hat{y}_{Norm,T}+\hat{y}_{Norm,C}$
4. Experiments
@@ 4.1. Dataset
  • Electricity (ecl, [n.d.]): This dataset contains electricity load diagrams for a power grid from 2011 to 2014, offering insights into energy consumption patterns over time. It presents a challenging forecasting task due to its seasonal variations and potential anomalies.
  • ETT (4 subsets) (Li et al., 2021): This dataset comprises four subsets with varying lengths and complexities, representing diverse time series forecasting challenges. It allows us to assess TEFN’s ability to handle different data characteristics and prediction horizons.
  • Exchange (Wu et al., 2021): This dataset captures exchange rate fluctuations between major currencies. It presents a dynamic and volatile environment, testing TEFN’s ability to capture complex market trends and predict future exchange rates accurately.
  • Traffic (tra, [n.d.]): This dataset contains traffic flow data collected from sensors in a city. It reflects the temporal and spatial patterns of traffic congestion, providing a challenging task for forecasting traffic flow and congestion levels.
  • Weather (wea, [n.d.]): This dataset includes historical weather data, encompassing various meteorological parameters. It allows us to evaluate TEFN’s performance in predicting weather patterns, which are crucial for numerous applications like agriculture, disaster management, and transportation planning.
Each of these datasets is characterized by its large size and diverse features, offering a rich and challenging environment for evaluating TEFN’s capabilities. The extensive nature of the data ensures that the model is tested on a wide range of scenarios, providing valuable insights into its performance and generalizability. More dataset features are in the Table 1.
Table 1. The quantitative features of the dataset
@@
Table 2. This table presents a comprehensive comparison of prediction errors across various models for long-term time series forecasting tasks. The models are evaluated on four different prediction lengths: $L_{pred}\in\{96,192,336,720\}$, providing a thorough assessment of their performance under varying forecasting horizons. The color coding in the table highlights the ranking of each model’s performance across all datasets and prediction lengths, allowing for easy identification of the top-performing models.
@@
9
0
0
1
1
0
3
0
4
0
0
0
0
0
$3^{\text{rd}}$ Count
8 5 9 9
5
6
12
11
3
0
0
0
6 11
0
1
1
0
3
0
4
0
0
0
0
0
Table 3. Full results of hyperparameter sensitivity
$|S|$ $lr$ Target Dataset
Electricity ETTh1 ETTh2 ETTm1 ETTm2 Exchange Traffic Weather
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
0 $1*10^{-2}$ 96 0.197 0.273 0.403 0.407 0.288 0.337 0.344 0.367 0.182 0.264 0.086 0.204 0.645 0.384 0.186 0.230
192 0.197 0.276 0.433 0.420 0.375 0.392 0.385 0.385 0.247 0.305 0.178 0.299 0.598 0.360 0.232 0.266
336 0.212 0.292 0.478 0.443 0.423 0.434 0.414 0.404 0.310 0.344 0.328 0.412 0.605 0.362 0.283 0.301
720 0.254 0.326 0.476 0.465 0.452 0.458 0.478 0.438 0.410 0.400 0.886 0.711 0.644 0.382 0.357 0.348
$5*10^{-2}$ 96 0.197 0.274 0.383 0.391 0.289 0.338 0.344 0.368 0.182 0.265 0.082 0.198 0.645 0.384 0.182 0.227
192 0.197 0.277 0.433 0.420 0.376 0.393 0.383 0.385 0.247 0.305 0.179 0.300 0.598 0.360 0.232 0.266
336 0.213 0.294 0.476 0.441 0.425 0.437 0.415 0.406 0.308 0.344 0.340 0.421 0.605 0.363 0.284 0.301
720 0.254 0.327 0.478 0.465 0.453 0.459 0.478 0.439 0.408 0.399 0.911 0.722 0.644 0.382 0.352 0.344
$1*10^{-1}$ 96 0.197 0.274 0.383 0.391 0.291 0.340 0.347 0.369 0.182 0.264 0.082 0.198 0.645 0.383 0.183 0.228
192 0.197 0.277 0.433 0.419 0.376 0.394 0.383 0.385 0.257 0.314 0.180 0.300 0.598 0.360 0.244 0.274
336 0.214 0.295 0.476 0.441 0.424 0.435 0.416 0.406 0.309 0.344 0.340 0.421 0.605 0.363 0.292 0.308
720 0.255 0.327 0.482 0.469 0.454 0.460 0.477 0.439 0.408 0.400 0.937 0.732 0.644 0.382 0.352 0.345
1 $1*10^{-2}$ 96 0.197 0.274 0.384 0.392 0.289 0.337 0.349 0.370 0.182 0.265 0.083 0.199 0.645 0.385 0.186 0.229
192 0.197 0.277 0.438 0.421 0.377 0.392 0.382 0.386 0.246 0.304 0.178 0.299 0.598 0.361 0.232 0.265
336 0.212 0.292 0.475 0.441 0.425 0.435 0.418 0.407 0.307 0.343 0.335 0.418 0.606 0.363 0.280 0.299
720 0.253 0.325 0.476 0.464 0.452 0.458 0.482 0.442 0.409 0.400 0.883 0.709 0.644 0.382 0.355 0.347
$5*10^{-2}$ 96 0.198 0.275 0.386 0.393 0.294 0.341 0.346 0.369 0.182 0.265 0.088 0.205 0.645 0.385 0.194 0.239
192 0.197 0.277 0.438 0.422 0.375 0.392 0.383 0.385 0.247 0.305 0.184 0.304 0.598 0.360 0.241 0.274
336 0.212 0.292 0.475 0.441 0.453 0.448 0.414 0.404 0.308 0.343 0.361 0.434 0.605 0.363 0.282 0.300
720 0.255 0.327 0.475 0.465 0.458 0.461 0.478 0.439 0.407 0.398 0.886 0.708 0.644 0.383 0.352 0.345
$1*10^{-1}$ 96 0.198 0.275 0.388 0.395 0.295 0.342 0.345 0.369 0.182 0.265 0.089 0.206 0.645 0.383 0.202 0.249
192 0.197 0.277 0.434 0.420 0.376 0.394 0.381 0.384 0.247 0.305 0.176 0.297 0.598 0.360 0.239 0.276
336 0.212 0.292 0.476 0.441 0.423 0.435 0.415 0.404 0.309 0.343 0.343 0.423 0.605 0.363 0.289 0.306
720 0.256 0.328 0.476 0.466 0.463 0.465 0.478 0.439 0.408 0.398 0.892 0.710 0.644 0.383 0.360 0.353
2 $1*10^{-2}$ 96 0.197 0.274 0.383 0.391 0.290 0.338 0.347 0.369 0.183 0.266 0.083 0.200 0.645 0.383 0.187 0.230
192 0.197 0.276 0.434 0.420 0.375 0.393 0.386 0.386 0.247 0.305 0.177 0.298 0.598 0.360 0.232 0.266
336 0.212 0.292 0.476 0.441 0.423 0.435 0.414 0.405 0.311 0.346 0.348 0.427 0.605 0.362 0.284 0.301
720 0.253 0.325 0.476 0.466 0.453 0.459 0.478 0.439 0.408 0.399 0.898 0.719 0.644 0.382 0.355 0.347
$5*10^{-2}$ 96 0.197 0.274 0.383 0.391 0.289 0.338 0.345 0.368 0.182 0.265 0.084 0.202 0.646 0.384 0.183 0.228
192 0.197 0.277 0.434 0.420 0.377 0.394 0.384 0.387 0.247 0.305 0.177 0.298 0.598 0.360 0.228 0.263
336 0.212 0.292 0.477 0.443 0.424 0.436 0.415 0.405 0.308 0.343 0.353 0.430 0.605 0.363 0.293 0.309
720 0.254 0.325 0.478 0.466 0.458 0.461 0.477 0.439 0.408 0.400 0.930 0.732 0.644 0.382 0.355 0.347
$1*10^{-1}$ 96 0.197 0.274 0.384 0.391 0.289 0.339 0.347 0.368 0.182 0.265 0.088 0.205 0.646 0.385 0.186 0.231
192 0.197 0.277 0.438 0.423 0.378 0.395 0.387 0.387 0.247 0.305 0.191 0.310 0.598 0.361 0.241 0.278
336 0.212 0.293 0.480 0.446 0.425 0.437 0.416 0.406 0.308 0.343 0.403 0.463 0.605 0.363 0.299 0.316
720 0.254 0.326 0.475 0.466 0.463 0.465 0.478 0.439 0.410 0.402 0.880 0.707 0.644 0.382 0.357 0.350
3 $1*10^{-2}$ 96 0.197 0.274 0.385 0.393 0.290 0.339 0.343 0.368 0.182 0.265 0.083 0.199 0.645 0.383 0.183 0.227
192 0.197 0.277 0.434 0.420 0.376 0.393 0.383 0.385 0.248 0.306 0.183 0.304 0.598 0.360 0.232 0.266
336 0.212 0.293 0.477 0.443 0.426 0.435 0.414 0.405 0.308 0.344 0.338 0.420 0.606 0.363 0.282 0.300
720 0.253 0.325 0.477 0.465 0.475 0.473 0.476 0.438 0.407 0.399 0.943 0.739 0.644 0.382 0.353 0.346
$5*10^{-2}$ 96 0.197 0.274 0.383 0.391 0.290 0.337 0.344 0.368 0.183 0.266 0.083 0.200 0.645 0.383 0.189 0.236
192 0.198 0.277 0.435 0.421 0.382 0.398 0.384 0.385 0.258 0.316 0.200 0.317 0.598 0.360 0.234 0.269
336 0.213 0.293 0.476 0.441 0.426 0.437 0.415 0.406 0.308 0.344 0.353 0.429 0.606 0.363 0.285 0.303
720 0.254 0.325 0.489 0.474 0.460 0.463 0.478 0.439 0.412 0.402 1.031 0.775 0.644 0.383 0.357 0.350
$1*10^{-1}$ 96 0.197 0.274 0.384 0.391 0.291 0.339 0.346 0.369 0.184 0.267 0.084 0.201 0.645 0.384 0.196 0.245
192 0.198 0.277 0.437 0.423 0.381 0.398 0.385 0.386 0.248 0.306 0.176 0.297 0.598 0.360 0.254 0.290
336 0.212 0.292 0.475 0.441 0.430 0.440 0.416 0.406 0.328 0.361 0.362 0.435 0.607 0.363 0.288 0.308
720 0.254 0.325 0.492 0.477 0.434 0.446 0.478 0.440 0.418 0.405 1.046 0.782 0.644 0.383 0.357 0.350
4 $1*10^{-2}$ 96 0.197 0.274 0.393 0.399 0.288 0.338 0.344 0.368 0.185 0.268 0.084 0.201 0.645 0.383 0.182 0.227
192 0.197 0.276 0.434 0.420 0.376 0.393 0.381 0.385 0.247 0.305 0.179 0.300 0.598 0.360 0.227 0.262
336 0.212 0.292 0.478 0.444 0.424 0.436 0.415 0.405 0.307 0.343 0.338 0.420 0.606 0.363 0.279 0.298
720 0.255 0.326 0.476 0.465 0.457 0.462 0.477 0.439 0.409 0.399 0.889 0.713 0.644 0.383 0.353 0.345
$5*10^{-2}$ 96 0.197 0.274 0.383 0.392 0.292 0.340 0.345 0.368 0.181 0.264 0.085 0.203 0.645 0.384 0.203 0.254
192 0.198 0.278 0.433 0.420 0.377 0.394 0.385 0.386 0.248 0.307 0.179 0.300 0.598 0.360 0.234 0.270
336 0.213 0.293 0.486 0.449 0.425 0.435 0.417 0.407 0.312 0.347 0.346 0.427 0.607 0.363 0.294 0.314
720 0.257 0.330 0.480 0.467 0.462 0.464 0.479 0.440 0.408 0.399 0.861 0.700 0.644 0.383 0.359 0.353
$1*10^{-1}$ 96 0.198 0.274 0.384 0.392 0.293 0.342 0.358 0.376 0.183 0.266 0.085 0.203 0.646 0.384 0.211 0.262
192 0.197 0.277 0.434 0.421 0.397 0.411 0.399 0.395 0.249 0.307 0.180 0.300 0.598 0.360 0.255 0.292
336 0.215 0.295 0.489 0.452 0.473 0.467 0.420 0.407 0.309 0.344 0.377 0.447 0.606 0.363 0.305 0.324
720 0.255 0.328 0.485 0.470 0.454 0.460 0.480 0.441 0.414 0.404 0.923 0.727 0.644 0.383 0.375 0.364
5 $1*10^{-2}$ 96 0.198 0.275 0.384 0.392 0.291 0.341 0.343 0.367 0.182 0.265 0.083 0.201 0.646 0.385 0.184 0.229
192 0.198 0.277 0.434 0.421 0.379 0.396 0.381 0.385 0.247 0.305 0.187 0.306 0.600 0.362 0.227 0.262
336 0.212 0.293 0.476 0.442 0.434 0.443 0.416 0.406 0.308 0.343 0.340 0.421 0.606 0.364 0.280 0.299
720 0.254 0.326 0.479 0.466 0.457 0.462 0.479 0.440 0.409 0.401 0.957 0.744 0.645 0.384 0.354 0.348
$5*10^{-2}$ 96 0.198 0.274 0.383 0.391 0.300 0.348 0.349 0.372 0.183 0.267 0.085 0.204 0.645 0.386 0.204 0.249
192 0.198 0.278 0.433 0.419 0.378 0.395 0.382 0.383 0.248 0.306 0.177 0.299 0.599 0.362 0.251 0.288
336 0.212 0.292 0.475 0.441 0.428 0.439 0.418 0.407 0.313 0.348 0.363 0.437 0.607 0.365 0.301 0.320
720 0.265 0.339 0.478 0.467 0.504 0.488 0.478 0.441 0.409 0.400 0.986 0.755 0.645 0.384 0.357 0.351
$1*10^{-1}$ 96 0.198 0.275 0.386 0.393 0.316 0.360 0.375 0.388 0.197 0.280 0.084 0.201 0.647 0.388 0.209 0.260
192 0.199 0.279 0.434 0.420 0.418 0.425 0.393 0.390 0.251 0.310 0.178 0.299 0.600 0.362 0.280 0.307
336 0.241 0.330 0.476 0.442 0.454 0.457 0.435 0.417 0.326 0.359 0.541 0.541 0.606 0.363 0.293 0.314
720 0.278 0.355 0.479 0.469 0.505 0.494 0.505 0.456 0.421 0.410 1.052 0.780 0.645 0.385 0.375 0.366
6 $1*10^{-2}$ 96 0.199 0.276 0.384 0.392 0.290 0.340 0.347 0.369 0.183 0.265 0.095 0.214 0.649 0.389 0.194 0.240
192 0.199 0.279 0.437 0.422 0.379 0.396 0.383 0.385 0.247 0.305 0.191 0.311 0.601 0.365 0.230 0.266
336 0.213 0.294 0.478 0.444 0.430 0.440 0.417 0.406 0.308 0.343 0.372 0.443 0.609 0.367 0.285 0.303
720 0.257 0.329 0.481 0.466 0.469 0.468 0.475 0.438 0.411 0.402 0.915 0.724 0.648 0.387 0.355 0.348
$5*10^{-2}$ 96 0.198 0.274 0.383 0.391 0.298 0.345 0.379 0.387 0.185 0.268 0.094 0.215 0.648 0.386 0.204 0.250
192 0.198 0.277 0.436 0.421 0.416 0.422 0.388 0.386 0.248 0.307 0.181 0.303 0.599 0.361 0.252 0.289
336 0.213 0.294 0.475 0.442 0.433 0.439 0.426 0.412 0.310 0.346 0.439 0.487 0.606 0.363 0.288 0.308
720 0.254 0.326 0.487 0.473 0.512 0.491 0.482 0.441 0.414 0.405 0.935 0.733 0.645 0.384 0.360 0.355
$1*10^{-1}$ 96 0.198 0.275 0.384 0.392 0.341 0.380 0.396 0.401 0.223 0.305 0.088 0.206 0.647 0.386 0.214 0.263
192 0.234 0.320 0.439 0.423 0.427 0.430 0.478 0.441 0.260 0.319 0.182 0.304 0.599 0.361 0.253 0.289
336 0.258 0.335 0.481 0.444 0.435 0.442 0.507 0.459 0.325 0.359 0.372 0.445 0.607 0.364 0.307 0.325
720 0.255 0.326 0.479 0.467 0.473 0.479 0.507 0.457 0.427 0.414 1.055 0.783 0.645 0.384 0.376 0.368
4.2. Comparative Baseline

### v3

@@ 3.3 Basic Probability Assignment Module
To introduce non-linear characteristics into TEFN and enhance its modeling capabilities, we can modify the fuzzy membership function used in the BPA module. Instead of the linear form presented in Equation (10), we can adopt non-linear functions that capture more complex relationships within the data. Several options exist for non-linear membership functions, including:
  • Gaussian Function: This function models the data distribution as a bell-shaped curve such as Equation (3), capturing the likelihood of each data point belonging to different membership levels. The parameters of the Gaussian function can be learned to adapt to the specific characteristics of the time series data.
  • Sigmoid Function: This function smoothly transitions between 0 and 1, providing a flexible way to represent the membership of data points to different levels in Equation (11). The sigmoid function can capture both gradual and abrupt changes in membership values.
    $\sigma(x)=\begin{cases}0&\text{if }x<0\\ \frac{1}{1+e^{-x}}&\text{if }x\geq 0\end{cases}$ (11)
  • Piecewise Functions: These functions consist of multiple linear or non-linear segments, allowing for more precise control over the shape of the membership function in Equation (12) where $g(x)$ is a non-linear function. This approach is useful for capturing complex patterns and relationships within the data.
    $f(x)=\begin{cases}a_{1}x+b_{1}&\text{if }c_{1}x+d_{1}\leq 0\\ g_{1}(x)&\text{if }c_{2}x+d_{2}>0\text{ and }c_{3}x+d_{3}\leq 0\\ a_{2}x+b_{2}&\text{if }c_{4}x+d_{4}>0\\ \vdots&\\ a_{n}x+b_{n}&\text{otherwise}\end{cases}$ (12)
@@
3.4 Expectation Fusion
TEFN addresses the challenge of handling multivariate time series data with two distinct dimensions: the time dimension $T$ and the channel dimension $C$. This results in the paralleled generation of different mass distributions for each dimension, denoted as $m_{C}$ and $m_{T}$. To integrate these mass distributions and produce a unified prediction, TEFN employs a fusion method based on expectations, as defined in Equation (13). Here, $y_{j,k}$ represents the fusion parameters, and $m^{\prime}_{i,j,k}$ corresponds to the transformed mass function for each dimension.
$\begin{split}\hat{y}_{Norm,D,i,j}=E_{D,F}(y_{j,k})&=\sum_{\begin{subarray}{c}j\in[1,|D|]\\ k\in[1,|2^{S}|]\end{subarray}}y_{j,k}*m^{\prime}_{i,j,k}\\ &=\sum_{\begin{subarray}{c}j\in[1,|D|]\\ k\in[1,|2^{S}|]\end{subarray}}m_{i,j,k}\end{split}$ (13)
The fusion operation in Equation (13) can be interpreted as a linear transformation of the mass functions, where the summation of $y_{j,k}*m^{\prime}_{i,j,k}$ effectively combines the information from different dimensions. The expectation is then calculated by summing the transformed mass functions, resulting in a single mass function representing the fused prediction.
TEFN deliberately avoids using the DSR for several reasons:
  1. 1.
    The computational complexity of DSR is significantly higher compared to the expectation-based fusion method. Additionally, DSR involves point-to-point multiplication of mass distributions within the same dimension, which can lead to accuracy loss when dealing with high-dimensional data due to finite precision limitations.
  2. 2.
    DSR is sensitive to extreme distributions, particularly those with a single element having a value of 1. This sensitivity can result in the final fused result being dominated by that single distribution, potentially leading to conflicts and loss of information from other distributions.
By employing the expectation-based fusion method, TEFN achieves a balance between computational efficiency and accuracy. This approach effectively integrates information from different dimensions while mitigating the risks of conflicts and accuracy loss associated with DSR. The final normalized prediction results are then aggregated across dimensions using the equation in Equation (14) and restored to their original scale through the de-normalization layer.
$\hat{y}_{norm}=\sum_{D\in\{T,C\}}\hat{y}_{Norm,D}=\hat{y}_{Norm,T}+\hat{y}_{Norm,C}$ (14)
4 Experiments
@@ 4.1 Dataset
  • Electricity [44]: This dataset contains electricity load diagrams for a power grid from 2011 to 2014, offering insights into energy consumption patterns over time. It presents a challenging forecasting task due to its seasonal variations and potential anomalies.
  • ETT (4 subsets) [19]: This dataset comprises four subsets with varying lengths and complexities, representing diverse time series forecasting challenges. It allows us to assess TEFN’s ability to handle different data characteristics and prediction horizons.
  • Exchange [45]: This dataset captures exchange rate fluctuations between major currencies. It presents a dynamic and volatile environment, testing TEFN’s ability to capture complex market trends and predict future exchange rates accurately.
  • Traffic [46]: This dataset contains traffic flow data collected from sensors in a city. It reflects the temporal and spatial patterns of traffic congestion, providing a challenging task for forecasting traffic flow and congestion levels.
  • Weather [47]: This dataset includes historical weather data, encompassing various meteorological parameters. It allows us to evaluate TEFN’s performance in predicting weather patterns, which are crucial for numerous applications like agriculture, disaster management, and transportation planning.
Each of these datasets is characterized by its large size and diverse features, offering a rich and challenging environment for evaluating TEFN’s capabilities. The extensive nature of the data ensures that the model is tested on a wide range of scenarios, providing valuable insights into its performance and generalizability. More dataset features are in the Table I.
TABLE I: The quantitative features of the dataset
@@
TABLE II: This table presents a comprehensive comparison of prediction errors across various models for long-term time series forecasting tasks. The models are evaluated on four different prediction lengths: $L_{pred}\in\{96,192,336,720\}$, providing a thorough assessment of their performance under varying forecasting horizons. The color coding in the table highlights the ranking of each model’s performance across all datasets and prediction lengths, allowing for easy identification of the top-performing models.
@@
4.2 Comparative Baseline